The length of a certain rectangle is 3 cm greater than its width. If the width were doubled and the length were decreased by 1 cm, the new rectangle would have the same perimeter as the original. Find the dimensions of the original rectangle.
Answer by edjones(8007)

L=w+3
P=2(w+w+3)=4w+6 replace L with w+3
.
2nd rectangle:
2w
L-1=w+2
2(2w+w+2)=4w+6
6w+4=4w+6
2w=2
w=1 cm, L=4 cm
.
Ed